OpenStudy (anonymous):

A bag countains 6 green marbles and 5 white select one at random what are the odds picking a green one?

OpenStudy (anonymous):

We have a total of 11 marbles (5+6=11) and there are 6 "correct" green ones. To find the probability we do: \[P(green)=\frac{correct}{total}=\frac{6}{11}\] which is also 0.5454 or about 54.4%

OpenStudy (anonymous):

thanks